Jen is three times as old as her son. Ten years from now, she will be twice as
old as her son will be then. How old is Jen’s son now?
I know J=s3
And I want to say J+10=2s+10, but that doesnt work.

Answer by rothauserc(4718)

J = 3s
J+10 = 2(s +10)
substitute for J in second equation
3s+10 = 2s +20
s = 10
J = 30
s = 10, Jenn's son is 10 years old